OpenStudy (anonymous):

george completed 3/5th of work in 9 days with the help of paul he completes remaining work in 4 days.In how many days paul alone can finish the work

OpenStudy (anonymous):

plzzz any one help me

OpenStudy (anonymous):

so the 1-3/5 =2/5 of the work was finished in 4 days or in other words Paul and Geroge finished 2/5 of the wor that is to say that Paul and Geroge together in 1 day can finish 2/(5*4) =1/10 of the work again Geroge alone in 1 day can do 3/(5*9) =1/15 of the work

OpenStudy (anonymous):

so in 1 day paul alone can do 1/10 -1/15 =(3-2)/30 =1/30 that means if Paul would complete the whole work in 30 days if working alone

OpenStudy (anonymous):

in these types of question always try to find the work or part of work done in 1 day

OpenStudy (anonymous):

if Paul would have been asked to complete the remaining 2/5 of the worked he would have taken 2/5 *30 =12 days (if working alone on the remaining portion of work)

OpenStudy (anonymous):

thanku :)

OpenStudy (anonymous):

welcome dear...

OpenStudy (ranga):

nicely done @matricked

OpenStudy (anonymous):

@ranga thank you

hero (hero):

George did 3/5 of the job in 9 days Had he continued working alone, George would completed the remaining 2/5 of the job in 6 days So George could do the whole job in 15 days Knowing this, we can figure out how fast George and Paul would have completed the entire job while working together. Let G = the time it takes George to do the whole job alone P = the time it takes Paul to do the whole job alone t = the time it takes them to complete the whole job together Then the following formula can be applied: \[\frac{GP}{G + P} = t\] We know G = 15 so \[\frac{15P}{15 + P} = t\] Also, we know that for the given problem they only completed 2/5 of the job working together for 4 hours so \[\frac{2}{5}\times \frac{15P}{15 + P} = 4\] And \[\frac{15P}{15 + P} = 10\] So had they worked together the whole time, it would have taken them only 10 hours to do the job. Now solving for P 15P = 10(15 + P) 15P = 150 + 10P 15P - 10P = 150 5P = 150 P = 150/5 P = 30 So it would take Paul 30 hours to do it alone.

hero (hero):

@ranga @matricked

OpenStudy (ranga):

@Hero We are all getting the same answers. My preference: Fraction of Work # of Days In 1 Day Fraction of Work Done George 3/5 9 1/15 (3/5 divided by 9) George + Paul 2/5 4 1/10 (2/5 divided by 4) Paul 1 x 1/x Work done in 1 day by both = work done in one day by each. Therefore, 1/10 = 1/15 + 1/x 1/x = 1/10 - 1/15 = 5/150 = 1/30 x = 30

hero (hero):

@ranga, I was just posting an alternative method. That is all.